OpenStudy (simplixity):

An archer fires an arrow straight up into the air with a speed vo = 42 m/s. Neglect air resistance. 1. Find the maximum height h reached by the arrow, in meters. 2.Write an expression for the time the arrow is in the air until it returns to launch height in terms of known quantities.

OpenStudy (3mar):

You need a direct answer or explanation?

OpenStudy (3mar):

anyway , we use the first law of motion Vf=Vo-a.t as it is a deceleration, not acceleration then t=4.281sec

OpenStudy (simplixity):

@3mar would I use that to find the height? or

OpenStudy (3mar):

No sir the hieght has its own equation which is Vf^2=Vo^2-a.d

OpenStudy (3mar):

Got it?

OpenStudy (simplixity):

@3mar so what would acceleration and distance equal if they only provide initial velocity?

OpenStudy (3mar):

acceleration is the gravitational deceleration -9.81 m/s2 as it go straight up.

OpenStudy (3mar):

The height equation is in the form: \[d=v.t+0.5a.t^2\]

OpenStudy (3mar):

v is the final velocity t is the time for complete climb and a is -9.81 m/s2 as you know Just get the d.

OpenStudy (3mar):

Is it simple now?

OpenStudy (simplixity):

sorry! how would I find time?

OpenStudy (3mar):

from the relation states: Vf=Vo+a.t Vf=0 Vo=42 a=-9.81 Got it?

OpenStudy (3mar):

This is the climb up time. Do the same for falling.

OpenStudy (3mar):

Excuse me, I have to go now. Could you allow?